OpenStudy (anonymous):

3x+6y=9 2x-10y=13 If I wanted to solve by substitution, what would be the best variable to solve for and in what equation?

OpenStudy (paki):

Solve one of the equations for one of the variables... We note that is simplest to solve the second equation for y... so... 3x+2y=−2x+1=3y SUBSTITUTE y into first equation. 3x+2(−2x+1)=−2x+1=3y Now solve first equation for x. x=−2x+1=−1y To find y, substitute −1 for x into second equation. x=y=−1−2⋅(−1)+1 The solution is: x=-1 and y=3 You can check the solution using the above calculator.

OpenStudy (anonymous):

That is wrong.. Solution is x=4 and y=-0.5

OpenStudy (texaschic101):

solve for y Because the first equation, when divided by 3, leaves you with x + 2y = 9. and x = -2y + 9....this can be subbed in for x in the 2nd equation, so then you can find y

OpenStudy (anonymous):

solving by graphing is different from solving by substitution. to solve by substitution, solve one equation for y in terms of x or x in terms of y. Then, substitute the resulting equation in the other equation for the appropriate variable. This gives you one equation in one variable, which you can solve in the usual way. Paki's answer is good.

OpenStudy (texaschic101):

oops..x + 2y = 3...not 9...my bad

OpenStudy (anonymous):

3x+6y=9 2x-10y=13 I will now make x the same. 6x-12y=18 6x-30=39 You would now subtract the two values so it is: 42y=-21 y=-0.5 Then you put it back into x and you work out that x=4

OpenStudy (texaschic101):

2x - 10y = 13 2(-2y + 3) - 10y = 13 -4y + 6 - 10y = 13 -4y -10y = 13 - 6 -14y = 7 y = -1/2 or -.05 yep...MathsPro is correct